40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 17122     Entry time: Wed Aug 31 11:39:48 2022
Author: Yehonathan 
Type: Update 
Category: LSC 
Subject: Updated XARM noise budget 

{Radhika, Paco, Yehonathan}

For educational purposes we update the XARM noise budget and add the POX11 calibrated dark noise contribution (attachment).

Attachment 1: Screenshot_2022-08-31_11-38-46.png  74 kB  Uploaded Wed Aug 31 12:41:37 2022  | Hide | Hide all
Screenshot_2022-08-31_11-38-46.png
ELOG V3.1.3-